How many five card hands consisting of 2 kings and 3 aces can be dealt from a deck of 52 playing cards?
I have it set up as C(52,2)=52!/(2!50!) & C(52,3)=52!/(3!49!)

Answer by stanbon(75887)

How many five card hands consisting of 2 kings and 3 aces can be dealt from a deck of 52 playing cards?
---------------
# of ways to succeed:: 4C2*4C3 = 6*4 = 24
# of random hands:: 52C5
---
Probability = 24/52C5 = 0.00000923....
Cheers,
Stan H.
========